There is a 30% probability that the train will be at least 2 minutes late. 1. Bob takes the train five times in one week. What is the probability that the train will only arrive late once? 2. What is the probability that the train will be late more than once? 3. Suppose Bob takes the train 22 times every month. What is the probability of being late exactly twice?
Added by Robert W.
Step 1
Given that there is a 30% probability that the train will be at least 2 minutes late, the probability of the train being on time is 1 - 0.30 = 0.70. Show more…
Show all steps
Close
Your feedback will help us improve your experience
Emily Anderson and 58 other Intro Stats / AP Statistics educators are ready to help you.
Ask a new question
Labs
Want to see this concept in action?
Explore this concept interactively to see how it behaves as you change inputs.
Key Concepts
Recommended Videos
Bob and Jane will each arrive at the train station at random times between one and three o’clock. Bob will wait 10 minutes; Jane will wait 20 minutes. What is the probability they will be at the station at the same time?
Emily A.
Recommended Textbooks
Elementary Statistics a Step by Step Approach
The Practice of Statistics for AP
Introductory Statistics
Transcript
Watch the video solution with this free unlock.
EMAIL
PASSWORD